OpenStudy (anonymous):

A colony of bacteria is growing exponentially, doubling in size every 140 minutes. How many minutes will it take for the colony of bacteria to become five times its current size?

OpenStudy (amistre64):

need to define the rate

OpenStudy (amistre64):

ln(2)/t = r just as a notion

OpenStudy (amistre64):

A = Pe^rt 2P = Pe^rt 2 = e^rt ln(2) = rt ln(2)/t = r .................... A = Pe^(ln(2)/t) t' A = Pe^(ln(2)/t) t' 5P = Pe^(ln(2)/t) t' 5 = e^(ln(2)/t) t' ln(5) = (ln(2)/t) t' ln(5)/(ln(2)/t) = t' maybe depending on what they current size might be?

OpenStudy (amistre64):

\[t'=log_2(5^{140})\] is my guess

OpenStudy (amistre64):

325.07 minutes ... IF im right lol

OpenStudy (anonymous):

yes you're right. I just wanted to check my answer. Thank you!!

OpenStudy (amistre64):

yep